# HOTS3D: Hyper-Spherical Optimal Transport for Semantic Alignment of Text-to-3D Generation

### v1

@@ Abstract
Recent CLIP-guided 3D generation methods have achieved promising results but struggle with generating faithful 3D shapes that conform with input text due to the gap between text and image embeddings. To this end, this paper proposes HOTS3D which makes the first attempt to effectively bridge this gap by aligning text features to the image features with spherical optimal transport (SOT). However, in high-dimensional situations, solving the SOT remains a challenge. To obtain the SOT map for high-dimensional features obtained from CLIP encoding of two modalities, we mathematically formulate and derive the solution based on Villani’s theorem, which can directly align two hyper-sphere distributions without manifold exponential maps. Furthermore, we implement it by leveraging input convex neural networks (ICNNs) for the optimal Kantorovich potential. With the optimally mapped features, a diffusion-based generator and a Nerf-based decoder are subsequently utilized to transform them into 3D shapes. Extensive qualitative and qualitative comparisons with state-of-the-arts demonstrate the superiority of the proposed HOTS3D for 3D shape generation, especially on the consistency with text semantics. The code will be publicly available.
@@ 1 Introduction
Text-to-3D generation endeavors to create 3D assets that are coherent with the input texts, which has the potential to benefit a wide array of applications such as animations, games, architecture, and virtual reality. Despite the extensive needs, the production of premium 3D content often remains a daunting task. The traditional process of creating 3D assets requires a significant investment of time and effort, even when undertaken by professional designers [1]. This challenge has prompted the development of text-to-3D methods. Recently, with the development of large language models (LLM) and large-scale datasets, deep learning based text-to-3D models [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] have made remarkable progress and can generate diverse shapes from various text prompts. The existing text-to-3D methods are either CLIP-guided end-to-end methods or optimization-based methods. The former trains a mapper from the CLIP [14] image embeddings to the shape embeddings of a 3D shape generator and switches to the CLIP text embedding as the input at test phase [15, 16, 2, 17, 18, 8]. The latter iteratively optimizes the CLIP similarity loss between a text prompt and renders images of a 3D scene representation [5, 4, 19, 3, 20] or the score distillation sampling loss [6, 7, 9, 10] to achieve semantic alignment.
@@
CLIP-guided optimization-based methods typically exhibit superior text-shape consistency. However, they require re-optimizing model parameters for each text input, resulting in significant generation time costs. For instance, DreamFields [5] takes over 30 minutes to generate a single 3D shape from a given text. In contrast, the CLIP-guided end-to-end methods demonstrate notably faster generation speed with acceptable quality. Despite their efficiency, these methods often overlook the gaps between text and image features as illustrated in the upper sphere in Fig. 1. This usually leads to the mismatching problem and semantic inconsistency of generated 3D shapes, especially in the testing stage which treats the two equally. Besides, as remarked in Dream3D [20], CLIP-guided methods tend to generate unconstrained “adversarial contents” that achieve high CLIP scores but with poor text consistency. To narrow the domain gap, Dream3D iteratively optimizes a text-to-image diffusion model, achieving more coherent results than other baselines. Nevertheless, as a CLIP-guided optimization-based method, Dream3D inherently suffers from extended sampling duration, which is further exacerbated by an additional diffusion process.
To improve the semantic consistency of text-to-3D generation while maintaining efficiency, we consider introducing optimal transport (OT) to align semantics and image features. Increasing OT-based methods are successfully applied in image generative models to align prior distribution for high quality and diverse generation [21, 22, 23, 24, 25]. However, they typically operate in Euclidean space, while there exists the constraint of normalization on the hyper-sphere for CLIP-guided text and image features, i.e. spherical optimal transport (SOT). Furthermore, solutions of SOT are relatively less investigated. Especially for hyper-spheres, the implementation is significantly challenging due to the high dimension.
To this end, we formulate and derive the SOT map for high-dimensional semantic and image feature distribution alignment, leading to a novel hyper-sphere-OT-based end-to-end text-to-3D generative model, named HOTS3D. For the implementation of SOT, we approach SOT as a spherical Kantorovich dual problem and employ two input convex neural networks (ICNN) [26] to represent two convex potentials [27], whose gradients are pivotal in solving the SOT map. Once the solution is extrapolated through the learning process, optimally mapped features are decoded into 3D shapes. Our main contributions are:
  • We are the first to formulate semantic alignment as a hyper-sphere OT problem for end-to-end text-to-3D generation. With the observation of hyper-spherical distributions of CLIP features, we propose to optimally transport text features to image features on hyper-spheres to bridge the semantic gap.
@@
  • We perform extensive experiments and analysis, which verify the superiority of the proposed HOTS3D in terms of semantic consistency and quality (lower FID score) compared with state-of-the-art methods.
@@ 2.1 Text-to-3D content generation
With the development of CLIP [14] technology and the need for 3D content creation, text-to-3D generation has attracted considerable attention [6, 5, 7, 11, 28, 29, 30, 31]. Existing methods can be categorized CLIP-guided end-to-end generative models with pretrained [15, 16, 2, 17, 32, 18, 8, 33, 34, 35, 36, 37] or fine-tuned CLIP encoders [38, 39, 40], and CLIP-guided optimization based methods [5, 4, 41, 19, 3, 20, 6, 42, 7, 9, 10, 43]. End-to-end generative models attempt to learn the distribution of the entire dataset, so that they can process different text inputs and generate various 3D assets without any fine-tuning. Optimization-based methods achieve bimodal alignment by iteratively optimizing the similarity between input text and the rendered images of the generated 3D assets. For those methods, it is necessary to finetune the model for each text, and the generation process is time-consuming.
In this work, we focus on end-to-end generative models. Among those existing methods, Shap·E [18] is most relevant to this paper, which can generate diverse objects without relying on images as an intermediate representation and has gained popularity in the research community due to its impressive results. Recently, Diffusion-SDF [17] proposed an SDF autoencoder together with the voxelized Diffusion model to learn and generate representations for voxelized signed distance fields (SDFs) of 3D shapes. SDFusion [8] enables 3D shape completion, reconstruction, and generation using a diffusion-based encoder-decoder. These CLIP-guided text-to-3D content generation methods have yielded impressive results, but as remarked in Dream3d [20], they overlooked the gap between the text and image embedding spaces. Consequently, semantic alignment is the key to filling this gap and improving text-shape consistency.
@@ 2.2 Semantic alignment
Most text-to-3D generation methods are proposed based on the assumption that image features and text features correspond in the cross-modal embedding space, which is usually extracted by CLIP. Due to the issue of inconsistent text shapes in generation models, more and more researchers are attempting to align semantic and image features further to improve text-shape consistency. For example, Li et al. [44] introduced a contrastive loss to align the image and text representations, making it easier for the multimodal encoder to perform cross-modal learning. Cheng et al. [45] designed a semantic alignment module to fully explore the latent correspondence between images and text. Guo et al. [46] modeled the distribution over discrete motion tokens which enables non-deterministic production of variable lengths pose sequences for input texts. 3D-VisTA [47] utilized a pre-trained Transformer for point cloud and text alignment which eliminates the need for auxiliary losses and optimization tricks in question answering tasks. Nie [48] mined the semantic consistency by compacting each 3D shape and its nearest neighbors to enhance semantic alignment for unlabeled 3D model domains. Cao [49] developed a cross-modal alignment module to align feature spaces between 3D point cloud and image/text modalities. For consistent generation, Zhao [50] proposed to use one network to encode the 3D shapes into the shape latent space aligned to the image and text, and use another network to learn a probabilistic map from the image or text space to the latent shape space. Liu et al. [41, 38, 42] finetuned the text encoder to improve the text-shape consistency. Xu et al. [20] bridged the text and image modalities with a diffusion model. These methods either only fine-tune the model from the perspective of cosine similarity, or require complex and time-consuming diffusion models, which motivates us to explore more effective semantic alignment methods.
@@ 2.3 Euclidean and spherical optimal transport
The OT map transmits one distribution to another in the most cost-effective manner, which has been widely applied for distribution alignment in Euclidean space [51, 21, 22, 23, 24, 52, 53, 54, 25, 55, 56, 57]. Recently, several works have shown that using a prior on the hyper-sphere can improve the results [58, 59, 60, 61]. Contrary to Euclidean geometry, both the metric and the transportation costs under the spheres differ, leading to the traditional OT algorithms being ineffectual for spheres. Consequently, OT solutions for spheres or hyper-spheres have attracted community attention. Cui et al. [62] provided a 2D SOT algorithm from a geometric view to adjust the density of triangular facets for mesh parameterization. Later, Hamfeldt et al. [63, 64] proposed a PDE approach to numerically solving the 2D SOT map and constrains the solution gradient to obtain a stable scheme. Cohen et al. [65] designed a flow model based on the Riemannian exponential map to fit the 2D SOT. However, due to the challenge of calculating Jacobian determinants (JD) in high-dimension, stochastic trace estimates need to be used to approximate JD for distribution transformation. Bonet et al. [65] defined spherical sliced-Wasserstein (SSW) distances on Riemannian manifolds and utilized the Flow model [65] to fit the 2D SOT by minimizing the SSW loss. In high-dimensional cases, SSW was taken as a regularization term to guide the decoder to output a feature distribution aligning with the prior distribution, but the map fitted by the decoder is not an SOT map. For the hyper-sphere OT, Rezende et al. [66] designed a Hessian matrix to reformulate the Kantorovich potential and gave a threshold criterion of the matrix for the hyper-sphere. Nonetheless, due to the difficulty in calculating the JD, only 2D SOT validation experiments were provided. As remarked in [66], solving the SOT problem in high-dimensional settings based on exact likelihoods is still challenging. Moreover, SOT algorithms like Ref [66, 65] used Riemannian exponential maps to switch Euclidean data to hyper-spheres and then calculated the OT map in a Euclidean way, which inevitably leads to distortion of manifolds and accuracy decline of results.
@@ 3 Method
Our objective is to generate 3D content that aligns with the given input text prompt. As illustrated in Fig. 2, our framework for text-guided 3D synthesis comprises four stages. Firstly, we encode the input text prompt onto the hyper-sphere with a pre-trained CLIP text encoder, obtaining text features. Secondly, the SOT map in Eq.(15) is induced by the gradient of a convex function (see Fig. 3) that is trained via minimax optimization, and then transfers output text features to the image feature space. In the third stage, a generator conditioned on the output of the SOT Map was utilized to generate 3D shape priors. This process involves a diffusion process based on a cross-attention score network. Finally, a Nerf-based decoder was utilized to decode the generated shape priors into 3D shapes. The SOT map is a plug-and-play tool for aligning spherical distributions. During the training phase, we only need to optimize the parameters of the SOT map and other modules remain frozen, significantly reducing the training difficulty.
Observations and Motivations. Though CLIP-guided end-to-end generative methods can generate diverse results, they are short in producing precise and detailed 3D structures that accurately match the text, as depicted in Fig. 5. Its cause lies in the gaps between text and image features, especially in the testing stage, treating the two equally, directly leading to a mismatch issue. To narrow this gap, CLIP-guided optimization methods attempt to align the semantics of instance text through iteration, resulting in significant time costs. Thus, from the perspective of distribution alignment, we propose using SOT Map to transfer entire text features into the image space. For solutions of SOT, existing works are either not applicable to hyper-sphere, or cause distortion, which encourages us to design a more effective method.
@@ 3.1 Problem formulation and preliminary
Given text feature set $\boldsymbol{X}\subset\mathbb{S}^{d}$, image feature $\boldsymbol{Y}\subset\mathbb{S}^{d}$, and their respective prescribed probability measure $\mu_{S}$ and $\nu_{S}$, our goal is to transfer text features to the image feature space to achieve alignment of two modalities, thereby providing more accurate features encoding for generations. In the following, we introduce relevant OT theories to provide preliminary for subsequent method explanations.
@@
Where $T$ achieving the minimum total cost in Eq. (1) is called OT map. $T_{\#}\mu$ is the push-forward of $\mu$ under $T$. Subsequently, by combining the joint distribution, Kantorovich relaxed the original Monge’s OT problem and provided a dual formulation Kantorovich duality problem [27],

### v2

# Hyper-spherical Optimal Transport for Semantic Alignment in Text-to-3D End-to-end Generation

@@ Abstract
Recent CLIP-guided 3D generation methods have achieved promising results but struggle with generating faithful 3D shapes that conform with input text due to the gap between text and image embeddings. To this end, this paper proposes HOTS3D which makes the first attempt to effectively bridge this gap by aligning text features to the image features with spherical optimal transport (SOT). However, in high-dimensional situations, solving the SOT remains a challenge. To obtain the SOT map for high-dimensional features obtained from CLIP encoding of two modalities, we mathematically formulate and derive the solution based on Villani’s theorem, which can directly align two hyper-sphere distributions without manifold exponential maps. Furthermore, we implement it by leveraging input convex neural networks (ICNNs) for the optimal Kantorovich potential. With the optimally mapped features, a diffusion-based generator is utilized to decode them into 3D shapes. Extensive quantitative and qualitative comparisons with state-of-the-art methods demonstrate the superiority of HOTS3D for text-to-3D generation, especially in the consistency with text semantics. Codes are available on HOTS3D.
@@ 1 Introduction
Text-to-3D generation endeavors to create 3D assets that are coherent with the input texts, which has the potential to benefit a wide array of applications such as animations, games, architecture, and virtual reality. Despite the extensive needs, the production of premium 3D content often remains a daunting task. Traditional process of creating 3D assets requires a significant investment of time and effort, even when undertaken by professional designers [1]. This challenge has prompted the development of text-to-3D methods. Recently, with the development of large language models and large-scale datasets, deep learning based text-to-3D models [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] have made remarkable progress and can generate diverse shapes from text prompts. The existing text-to-3D methods are either CLIP-guided end-to-end methods or optimization-based methods. The former trains a mapper from CLIP [13] image embeddings to the shape embeddings of a 3D shape generator and switches to the CLIP text embedding as the input at test phase [14, 15, 16, 17, 18, 7]. The latter iteratively optimizes the CLIP similarity (SDS) loss between a text prompt and renders images of a 3D scene representation [4, 3, 19, 2, 20] or the score distillation sampling loss [5, 6, 8, 9] to achieve semantic alignment.
@@
CLIP-guided optimization-based methods typically exhibit superior text-shape consistency. However, they require re-optimizing model parameters for each text input, resulting in significant generation time costs. For instance, ProlificDreamer [9] takes over ten hours to generate a single 3D shape from a given text. In contrast, the end-to-end methods demonstrate notably faster generation speed with acceptable quality. Despite their efficiency, these methods often overlook the gaps between text and image features as illustrated in the upper sphere in Fig. 1. This usually leads to a mismatch and semantic inconsistency of generated 3D shapes, especially in the testing stage, which treats the two equally. Besides, as remarked in Dream3D [20], CLIP-guided methods tend to generate unconstrained “adversarial contents” that achieve high CLIP scores but with poor text consistency. To narrow the domain gap, Dream3D iteratively optimizes a text-to-image diffusion model, achieving more coherent results than other baselines. Nevertheless, as an optimization-based method, Dream3D inherently suffers from extended sampling duration, which is further exacerbated by an additional diffusion process.
To improve the semantic consistency of text-to-3D generation while maintaining efficiency, we consider introducing optimal transport (OT) to align semantics and image features. Increasing OT-based methods are successfully applied in image generative models to align prior distribution for diverse generation [21, 22, 23, 24, 25]. However, they typically operate in Euclidean space, while there exists the constraint of normalization on the hyper-sphere for CLIP-guided text and image features, i.e. spherical optimal transport (SOT). Furthermore, solutions for SOT are relatively less investigated. Especially for hyper-spheres, the implementation is challenging due to the high dimension.
To this end, we formulate and derive the SOT map for high-dimensional semantic and image feature distribution alignment, leading to a novel hyper-sphere-OT-based end-to-end text-to-3D generative model, named HOTS3D. For SOT implementation, we approach SOT as a spherical Kantorovich dual problem and employ two input convex neural networks (ICNN) [26] to represent two convex potentials [27], whose gradients are pivotal in solving SOT map. Once the solution is extrapolated, optimally mapped features are decoded into shapes. Our main contributions are:
  • We are the first to formulate semantic alignment as a hypersphere OT problem for end-to-end text-to-3D generation. With the hyper-spherical constraint of CLIP features, we propose to optimally transport text features to image features on hyper-spheres to bridge the semantic gap.
@@
  • We perform extensive experiments and analysis, which verify the superiority of the proposed HOTS3D in terms of semantic consistency and quality (lower FID, P-FID, and LFD score) compared with state-of-the-art methods.
Figure 2: Monster Invasion, generated by HOTS3D
@@ 2.1 Text-to-3D content generation
With the development of CLIP [13] technology and the need for 3D content creation with simple text prompts, text-to-3D generation has attracted considerable attention [5, 4, 6, 10, 28, 29, 30, 31]. Existing methods can be categorized CLIP-guided end-to-end generative models with pretrained [14, 15, 16, 17, 32, 18, 7, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42] or fine-tuned CLIP encoders [43, 44, 45], and CLIP-guided optimization based methods [4, 3, 46, 19, 2, 20, 5, 47, 6, 8, 9, 48, 49, 50, 51, 52, 48, 53, 54]. End-to-end generative models attempt to learn the distribution of the entire dataset, so that they can process different text inputs and generate various 3D assets without any fine-tuning. Optimization-based methods achieve bimodal alignment by iteratively optimizing the similarity between input text and the rendered images of the generated 3D assets. For those methods, it is necessary to finetune the model for each text, and the generation process is time-consuming. To facilitate the selection of baselines for experiments, we conducted a comprehensive investigation of the latest state-of-the-art baselines and summarized them in Tab. I.
In this work, we focus on more efficient text-guided generation models, i.e. end-to-end models. Among those existing methods, Shap·E [18] is most relevant to this paper, which can generate diverse objects without relying on images as an intermediate representation and has gained popularity in the research community due to its impressive results. Recently, Diffusion-SDF [17] proposed an SDF autoencoder together with the voxelized Diffusion model to learn and generate representations for voxelized signed distance fields (SDFs) of 3D shapes. SDFusion [7] enables 3D shape completion, reconstruction, and generation using a diffusion-based encoder-decoder. EXIM [39] adopts a two-stage coarse-to-fine pipeline to generate shapes without time-consuming per-shape optimization. However, it requires training separate models for each category, such as chairs and tables, and the corresponding model of the target category needs to be called during generation. Direct3D [41] focuses on image-to-3D generation, while text-to-3D generation is achieved by integrating existing text-to-image algorithms. Michelangelo [40] aligns shape and clip text/image features through autoencoders to generate high-quality shapes, but there are still significant semantic inconsistencies in publicly available models. These CLIP-guided text-to-3D content generation methods have yielded impressive results, but as remarked in Dream3d [20, 40], most of them overlooked the gap between the text and image embedding spaces. Consequently, semantic alignment is the key to filling this gap and improving text-shape consistency.
TABLE I: Recent works for text-to-3D generation (as of 12/2024). E2E, Open, and Size denote end-to-end, open source code and models, and the size of test prompts. SIGA, SPN, OBJ, 3DF, and NA denote SIGGRAPH Asia, ShapeNet, Objaverse, 3D-FUTURE, and ’means unknown’ respectively.
Methods Venue E2E Open 3D dataset Size
Kiss3DGen [55] arXiv 2025 OBJ 100
TRELLIS [56] arXiv 2024 OBJ NA
3DTopia-XL [57] arXiv 2024 OBJ NA
Ln3diff [58] ECCV 2024 SPN & OBJ NA
Xcube [38] CVPR 2024 SPN & OBJ 30
EXIM [39] TOG 2023 SPN & 3DF NA
Michelangelo [40] NIPS 2024 SPN 2592
Diffusion-SDF [17] CVPR 2023 Text2Shape NA
[…]
TAPS3D [33] CVPR 2023 SPN 1000
SDFusion [7] CVPR 2023 Text2shape NA
CLIP-Forge [16] CVPR 2022 SPN 234
Hunyuan3D [59] arXiv 2024 × OBJ 70
DreamMesh [54] ECCV 2024 × T3Bench 50
DiverseDream [60] ECCV 2024 × None 70
COMOGen [53] arXiv 2024 × OBJ 20
CLAY [42] TOG 2024 × SPN & OBJ 50
RichDreamer [48] CVPR 2024 × OBJ 113
LucidDreamer [52] CVPR 2024 × Point-e 28
VP3D [50] CVPR 2024 × T3Bench 300
GaussDreamer [49] CVPR 2024 × Point-e NA
ProlificDreamer [9] NIPS 2024 × None 100
3DFuse [51] ICLR 2024 × Co3D 42
Magic3D [6] CVPR 2023 × None 397
Dreamfusion [5] ICLR 2023 × None 153
PCLIPNeRF [2] arXiv 2022 × None 153
CLIP-Mesh [19] SIGA 2022 × None NA
@@ 2.2 Semantic alignment
Most text-to-3D generation methods are proposed based on the assumption that image features and text features correspond in the cross-modal embedding space, which is usually extracted by CLIP. Due to the issue of inconsistent text shapes in generation models, more and more researchers are attempting to align semantic and image features further to improve text-shape consistency. For example, Li et al. [61] introduced a contrastive loss to align the image and text representations, making it easier for the multimodal encoder to perform cross-modal learning. Cheng et al. [62] designed a semantic alignment module to fully explore the latent correspondence between images and text. Guo et al. [63] modeled the distribution over discrete motion tokens, which enables non-deterministic production of variable-length pose sequences for input texts. 3D-VisTA [64] utilized a pre-trained Transformer for point cloud and text alignment, eliminating the need for auxiliary losses and optimization tricks in question-answering tasks. Nie [65] mined the semantic consistency by compacting each 3D shape and its nearest neighbors to enhance semantic alignment for unlabeled 3D model domains. Cao [66] developed a cross-modal alignment module to align feature spaces between 3D point clouds and image/text modalities. For consistent generation, Zhao [40] proposed using one network to encode the 3D shapes into the shape latent space aligned to the image and text, and using another network to learn a probabilistic map from the image or text space to the latent shape space. Liu et al. [46, 43, 47] finetuned the text encoder to improve the text-shape consistency. Xu et al. [20] bridged the text and image modalities with a diffusion model. These methods require iterative optimization of models with cosine similarity, or complex and time-consuming diffusion models, which motivates us to explore more effective alignment methods.
@@ 2.3 Euclidean and spherical optimal transport
The OT map transmits one distribution to another in the most cost-effective manner, which has been widely applied for distribution alignment in Euclidean space [67, 21, 22, 23, 24, 68, 69, 70, 25, 71, 72, 73]. Recently, several works have shown that using a prior on the hyper-sphere can improve the results [74, 75, 76, 77]. Contrary to Euclidean geometry, both the metric and the transportation costs under the spheres differ, leading to the traditional OT algorithms being ineffectual for spheres. Consequently, OT solutions for spheres or hyper-spheres have attracted community attention. Cui et al. [78] provided a 2D SOT algorithm from a geometric view to adjust the density of triangular facets for mesh parameterization. Later, Hamfeldt et al. [79, 80] proposed a PDE approach to numerically solving the 2D SOT map and constraining the solution gradient to obtain a stable scheme. Cohen et al. [81] designed a flow model based on the Riemannian exponential map to fit the 2D SOT. However, due to the challenge of calculating Jacobian determinants (JD) in high dimensions, stochastic trace estimates need to be used to approximate JD for distribution transformation. Bonet et al. [77] defined spherical sliced Wasserstein (SSW) distances on Riemannian manifolds and utilized the Flow model [82] to fit the 2D SOT by minimizing the SSW loss. In high-dimensional cases, SSW was taken as a regularization term to guide the decoder to output a feature distribution aligning with the prior distribution, but the map fitted by the decoder is not an exact SOT map. For the hyper-sphere OT, Rezende et al. [83] designed a Hessian matrix to reformulate the Kantorovich potential and gave a threshold criterion of the matrix for the hyper-sphere. Nonetheless, due to the difficulty in calculating the JD, only 2D SOT validation experiments were provided. As remarked in [83], solving the SOT problem in high-dimensional settings based on exact likelihoods is still challenging. Moreover, SOT algorithms like Ref [83, 81] used Riemannian exponential maps to switch Euclidean data to hyper-spheres and then calculated the OT map in a Euclidean way, which inevitably leads to distortion of manifolds and accuracy decline of results.
@@ 3 Method
Our objective is to generate 3D content that aligns with the given input text prompt. As illustrated in Fig. 3, our framework for text-guided 3D synthesis comprises three stages. Firstly, we encode the input text prompt onto the hyper-sphere with a pre-trained CLIP text encoder, obtaining text features. Secondly, the SOT map in Eq.(15) is induced by the gradient of a convex function (see Fig. 4) that is trained via minimax optimization, and then transfers output text features to the image feature space. In the third stage, a generator conditioned on the output of the SOT Map was utilized to generate 3D shapes. The SOT map is a plug-and-play tool for aligning spherical distributions. During the training phase, we only need to optimize the parameters of the SOT map and other modules remain frozen, significantly reducing the training difficulty. With the SOT map for semantic alignment, our HOTS3D can bypass iterative optimization during the testing phase, resulting in stronger generalization capability and semantic consistency.
Observations and Motivations. Though CLIP-guided E2E methods can generate diverse results, they are short in producing precise and detailed 3D structures that accurately match the text, as depicted in Fig. 2 and Fig. 6. Its cause lies in the gaps between text and image features, especially in the testing stage, treating the two equally, directly leading to a mismatch issue. To narrow this gap, CLIP-guided optimization methods attempt to align the semantics of instance text through iteration, resulting in significant time costs. Thus, from the perspective of distribution alignment, we propose using SOT Map to transfer entire text features into the image space. For solutions of SOT, existing works are either not applicable to hyper-sphere, or cause distortion, which encourages us to design a more effective method.
@@ 3.1 Problem formulation and preliminary
Given text feature set $\boldsymbol{X}\subset\mathbb{S}^{d}$, image feature $\boldsymbol{Y}\subset\mathbb{S}^{d}$, and their respective prescribed probability measure $\mu_{S}$ and $\nu_{S}$, our goal is to transfer text features to the image feature space to achieve alignment of two modalities, thereby providing more accurate features encoding for generations. In the following, we introduce relevant OT theories to provide a preliminary for subsequent method explanations.
@@
$T$ achieving the minimum total cost in Eq. (1) is called OT map. $T_{\#}\mu$ is the push-forward of $\mu$ under $T$. Subsequently, by combining the joint distribution, Kantorovich relaxed the original Monge’s OT problem and provided a dual formulation Kantorovich duality problem [27],